OpenStudy (anonymous):

Can someone please help me? 5+(sqrt(x+2))=8+(sqrt(x-7))

myininaya (myininaya):

solving for x I assume

myininaya (myininaya):

I would isolate one of the sqrt( ) things then square both sides

myininaya (myininaya):

\[\sqrt{x+2}=3+\sqrt{x-7}\] square both sides what do you get after doing just this step

OpenStudy (anonymous):

Do I square both sides?

myininaya (myininaya):

yes that is what I'm asking you to do :p

OpenStudy (anonymous):

x+2=9+x-7?

myininaya (myininaya):

well that isn't exactly how you square the right hand side

OpenStudy (anonymous):

How would you do it then. That's the part I'm so confused on lol

myininaya (myininaya):

\[(\sqrt{x+2})^2=(3+\sqrt{x-7})^2 \\ x+2=(3+\sqrt{x-7})(3+\sqrt{x-7})\]

OpenStudy (anonymous):

Ok that would make sense..But then what..?

myininaya (myininaya):

multiply

OpenStudy (anonymous):

\[x+2=9+6(\sqrt{x-7})^2+x-7\]

OpenStudy (anonymous):

Is that right?

myininaya (myininaya):

well almost

myininaya (myininaya):

the square there shouldn't be on that radical part

myininaya (myininaya):

\[x+2=9+6 \sqrt{x-7}+x-7 \]

myininaya (myininaya):

now we want to isolate the square root item and then square both sides again

myininaya (myininaya):

subtract x on both sides + 7 on both sides - 9 on both sides

myininaya (myininaya):

\[7-9+2=6 \sqrt{x-7}\]

myininaya (myininaya):

combine like terms on left hand side

myininaya (myininaya):

divide both sides by 6

myininaya (myininaya):

and then square both sides

OpenStudy (anonymous):

You are my hero xD Thank you so much!!!

myininaya (myininaya):

make sure you check the answer by pluggin it into the original equation